# Hunting Primordial Black Hole Dark Matter in Lyman-$\alpha$ Forest

### v1

@@ Abstract
A very pressing question in contemporary physics is the identity of Dark Matter (DM), and one that has not been answered affirmatively to any degree so far. Primordial Black Holes (PBHs) are one of the most well-motivated DM candidates. Light enough PBHs have been constrained by either the non-detection of their Hawking radiation itself, or by the non-observation of any measurable effects of this radiation on astrophysical and cosmological observables. We constrain the PBH density by their Hawking radiation effect on the intergalactic medium (IGM) temperature evolution. We use the latest deductions of IGM temperature from Lyman-$\alpha$ forest observations. We put constraints on the fraction of PBH DM with masses $5\times 10^{15}$ g – $10^{17}$ g, separately for spinning and non-spinning BHs. We derive constraints by dealing with the heating effects of the astrophysical reionization of the IGM in two ways. In one way, we completely neglect this heating due to astrophysical sources, thus giving us weaker constraints, but completely robust to the reionization history of the universe. In the second way, we utilise some modelling of the ionization and temperature history, and use it to derive more stringent constraints. We find that for non-spinning PBHs of mass $10^{16}$ g, the current measurements can constrain the PBH-density to be $\lesssim$ 0.1% of the total DM. We find that these constraints from the latest Lyman-$\alpha$ forest temperature measurements are competitive, and hence provide a new observable to probe the nature of PBH DM. The systematics affecting Lyman-$\alpha$ forest measurements are different from other constraining observations, and thus this is a complementary probe.
@@ I Introduction
Several astrophysical and cosmological evidences point towards the existence of Dark Matter (DM), something that makes up around $\sim$ 85% of matter content today. But all these evidences rely on the gravitational interactions of DM. The non-gravitational nature of DM is yet to be discovered. Due to a wide variety of DM models and a huge range in DM candidate masses, there has been a plethora of research on the possibility and methods of observing non-gravitational imprints of DM [1, 2, 3, 4].
@@
Primordial Black Hole [5, 6, 7, 8, 9] (PBH) is one of the most popular and widely discussed DM candidates [10, 11, 12]. Various production mechanisms for PBH DM have been proposed in the literature and various aspects concerning their formation have been studied [13, 14, 15, 16, 17, 6, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Irrespective of its production mechanism, a PBH would emit Hawking radiation, the spectrum of which depends on the BH mass, charge, and spin [8, 35, 36, 37, 38, 39, 40]. Several terrestrial and space-based telescopes have been used to directly search for Hawking evaporation from PBH DM. These efforts have resulted in very stringent constraints on the abundance of low-mass ($\sim$ $10^{15}$ g - $5\times 10^{17}$ g ) PBH DM [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. Besides these direct searches for Hawking evaporation from PBH DM, a number of techniques have also been explored to search for indirect imprints of this radiation. Standard Model (SM) particles emitted by evaporating PBH DM can dump their energy into the surroundings. Effects of this exotic energy injection have been searched using various cosmological observables like Cosmic Microwave Background [54, 55, 56, 57] and global 21 cm signal [58, 59, 60, 61, 62, 63, 64].
@@
Lyman-$\alpha$ ($\sim$1216 Å) forest, which is a series of absorption lines seen in the spectra of distant quasars, remains one of the most important tracers of the diffuse inter-galactic medium (IGM) [65, 66, 67, 68, 69, 70]. Besides being sensitive to various cosmological parameters and reionizing sources [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86], observations of these Lyman-$\alpha$ absorption features are sensitive to the thermal and ionization history of the IGM [65, 66, 67, 68, 87, 88, 89, 90, 91, 92]. With the current and promising progress in Lyman-$\alpha$ forest observations, it is important to investigate the potential of these observations to uncover Beyond the Standard Model (BSM) physics. Several works have already demonstrated the power of Lyman-$\alpha$ forest in searching for various BSM scenarios [93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107]. In this work, Lyman-$\alpha$ forest observations are used for the first time to put limits on the abundance of low-mass evaporating PBH DM.
@@
The whole paper is structured as follows. Section (II) gives an overview of the energy injection into IGM by evaporating PBHs. In section (III), we briefly describe the evolution of the IGM temperture and ionization history in presence of Hawking evaporating PBHs. The first part of section (IV) is a brisk tour of the Lyman-$\alpha$ forest observations in the context of IGM temperature measurements. Then it lists the various datasets of temperature measurements using this technique that we make use of. We discuss our results in section (V) and then conclude in section (VI).
@@ II Energy injection by evaporating PBH DM
A PBH with a given mass and angular momentum will evaporate via Hawking radiation [8]. Temperature of an uncharged spinning PBH is,
@@
with $a_{*}$, $J$, and $M_{\rm PBH}$ being the reduced spin parameter, magnitude of angular momentum, and mass of the PBH, respectively, and $G$ is the gravitational constant. For a non-spinning PBH ($a_{*}=0$), the above expression simplifies to,
@@
During Hawking evaporation, the number of particles, $N_{i}$ emitted per unit energy per unit time is given by  [35, 36, 38, 39]
@@
where $\Gamma_{i}$ is the greybody factor  [35, 36, 38] encoding the probability of an emitted particle overcoming the gravitational potential of a black hole, and $E$ is the total energy of the emitted particles, $m$ is the axial quantum number, $s$ is the spin of the emitted particle, $\mu$ is the rest mass of the emitted particle, and $dof$ stands for the degrees of freedom associated with the emitted particles.
@@
Figure 2: Constraints on non-spinning ($a_{*}$ = 0) (left panel) and spinning ($a_{*}$ = 0.9999) (right panel) PBH DM (monochromatic mass distribution). Constraints from the conservative model are shown in red (deep red and the shaded region above that), whereas the photoheating model limits are shown in blue (deep blue and the shaded region above that). The deep blue and deep red bands are obtained by varying over the photoheating models used in this work (FlexKnot [4], Tanh [4], and Trac [108, 109]). Previous constraints include low-energy positron measurements from Galactic Centre by INTEGRAL (navy [43] and green [112]), cosmic-ray flux measurement by Voyager (purple) [112], INTEGRAL (orange) and COMPTEL (teal) measurements of the Galactic gamma-ray flux [46, 45], Super-Kamiokande limit for diffuse supernovae neutrino background (yellow) [47], measurement of extra-Galactic gamma-ray emission (aqua) [48, 49], gas heating in Leo T (black) [50, 51], PLANCK measurement of CMB (lime) [54], and XMM-NEWTON measurement of Galactic diffuse X-ray emission (pink) [112]. Some of the bounds for non-spinning PBH DM are absent for spinning PBH DM simply because they have not been derived in the literature.
In this work, the Hawking evaporation spectra are obtained using the publicly available code, BlackHawk v2.2[39, 113]. We have explicitly matched the spectra from BlackHawk with ref. [35]. At low energies, for secondary photon and electron-positron spectra, BlackHawk uses public Python code Hazma[114]. In this work we have used the total emission spectra (primary and secondary) from PBHs in the mass range  $\sim\,5\times 10^{15}\,{\rm g}-10^{17}$ g. Note that the only species of particles whose spectra we had to consider were photons and $e^{\pm}$. This is because only these particles are capable of efficiently interacting with the gas in the IGM and substantially affecting its thermal and ionization state [115, 116]. Other stable particles that are abundantly present in the Hawking radiation emitted by black holes of these masses are neutrinos and anti-neutrinos, and these have no effect on the temperature evolution of the IGM due to their weak-interaction.
@@
where $dV$ is the volume element and $f_{c}(z)$ is the deposition efficiency into channel $c$ at a given redshift $z$[118, 119, 120], respectively. The relevant channels are hydrogen ionization ($f_{ion}$), helium ionization ($f_{He}$), hydrogen excitation ($f_{exc}$), and heating of IGM ($f_{heat}$). The $f_{c}(z)$s have been computed in refs. [118, 119, 120] and have been implemented in DarkHistory[121, 100, 122, 123, 124, 125, 126]. In our work we have used DarkHistory to calculate the evolution of IGM temperature and ionization in presence of evaporating PBHs.
@@
×$\times\left(\frac{d^{2}N}{dEdt}\right)_{e^{\pm}}n_{\rm PBH}(M_{\rm PBH})\,\Psi(M_{\rm PBH})\,dM_{\rm PBH}\,dE_{e}.$ (9)
@@
with $f_{\rm PBH}$, $\rho_{c}$, and $\Omega_{\rm DM}$ being the fraction of DM in the form of PBHs, the critical density, and DM density parameter in the present universe. In our work we have used the monochromatic mass distribution of PBH DM, i.e., all PBHs are assumed to be of the same mass. Thus, $\Psi(M_{\rm PBH})$ takes the form of a dirac delta function peaked at PBH mass $M_{\rm PBH}$.
III Ionization and Thermal history of IGM
We describe the effects of evaporating PBHs on the IGM temperature and hydrogen fraction. The rate of change of IGM temperature ($T_{m}$) is given by [127, 128, 129, 121],
@@ III Ionization and Thermal history of IGM
In the above equations, $H$ is the Hubble parameter, $T_{\text{CMB}}$ is the Cosmic Microwave Background (CMB) temperature, $x_{e}$ is the free electron fraction ($\equiv n_{\text{e}}/{n_{\text{H}}}$). The parameter, $f_{\rm He}\equiv n_{\text{He}}/{n_{\text{H}}}$, where $n_{\text{He}}$ and $n_{\text{H}}$ are the number densities of helium and hydrogen nuclei, respectively; $\sigma_{T}$ is the Thomson cross section, $a_{r}$ is the radiation constant, and $m_{e}$ is the electron mass. The term $\dot{T}_{\text{atom}}$ in Eq. (12) represents the cooling term due to all atomic processes like collisional ionization, collisional excitation and Coulomb heating [118]. Finally, the $\dot{T}^{\star}$ term in Eq. (12) is the photoheating term that accounts for the heating due to photoionization from various astrophysical sources. This term depends on the ionization history of the IGM, which we briefly describe in what follows.
The evolution equation for IGM hydrogen fraction is:
@@
where $x_{\text{HII}}=n_{\text{HII}}/{n_{\text{H}}}$, with $n_{\text{HII}}$ being the number density of ionized hydrogen. The first term on the right hand side is again the contribution due to astrophysical sources and the second term represents the contribution due to PBH energy injection. The first term in Eq. (15) is:
@@
Here, $x_{\text{e}}^{\text{P}}$ is chosen from the Planck 95% confidence region limits on free electron fraction using ‘Tanh’ and ‘FlexKnot’ models [4]. In addition, we also use the 95% confidence region limit from parametrization described in Refs. [108, 109], hereafter referred to as ‘Trac’ model. We note that, at sufficiently high redshifts  [^1] Please refer to Ref. [100] for more details on how to fix redshift above which $\dot{x}_{\text{HII}}^{\star}=0$., $\dot{x}_{\text{HII}}^{\star}=0$.
Finally, we use the following prescription to fix $\dot{T}^{\star}$,
$\dot{T}^{\star}=\begin{dcases}\dot{x}_{\text{HII}}^{\star}(1+f_{\rm He})\Delta T\,,\,\,\,\,&x_{\text{HII}}<0.99\,,\\ \sum_{\text{X}=\{\text{H},\text{He}\}}\frac{h\nu_{\text{X}}}{3(\gamma_{\text{X}}-1+\alpha_{\text{bkg}})}\alpha_{\text{A},\text{X}}\,n_{\text{X}}\,,&x_{\text{HII}}\geq 0.99\,,\end{dcases}$ (18)
In the above equation, $\Delta T$ is a constant whose value lies between $2\times 10^{4}\,{\rm K}-3\times 10^{4}$ K as discussed in Refs. [133, 134, 88, 135, 92]. Here, $\nu_{\text{X}}$ is the frequency corresponding to the ionization potential of species $\text{X}\in\{\text{HI},\text{HeI}\}$, $h$ is Planck’s constant, $\gamma_{\text{X}}$ is the index for photoionization cross-section, $\alpha_{\text{bkg}}$ is the index for power-law ionizing background [134, 92, 91], and $\alpha_{\text{A},\text{X}}$ are the case-A recombination factors associated with the transitions XI $\rightarrow$ X [92].
@@
where $\mathcal{R}=13.6$ eV is the hydrogen ionization potential and $\mathcal{C}$ is the Peebles C factor [127, 129].

### v2

# Hunting Primordial Black Hole Dark Matter in the Lyman-$\alpha$ Forest

@@ Abstract
A very pressing question in contemporary physics is the identity of Dark Matter (DM). Primordial Black Holes (PBHs) are one of the most well-motivated DM candidates. Light PBHs have been constrained by either the non-detection of their Hawking radiation itself, or by the non-observation of any measurable effects of this radiation on astrophysical and cosmological observables. We constrain the PBH contribution to the DM density by non-detection of their Hawking radiation’s effect on the intergalactic medium (IGM) temperature evolution. We use the latest deductions of IGM temperature from Lyman-$\alpha$ forest observations. We put constraints on the fraction of DM as PBHs with masses $5\times 10^{15}$ g – $10^{17}$ g, separately for spinning and non-spinning BHs. We derive constraints by dealing with the heating effects of the astrophysical reionization sources on the IGM in two ways. In one way, we completely neglect this heating due to astrophysical sources, thus giving us weaker constraints, but completely robust to the reionization history of the universe. In the second way, we utilise some modelling of the ionization and temperature history, and use it to derive more stringent constraints. We find that for non-spinning PBHs of mass $10^{16}$ g, the current measurements can constrain the PBH-density to be $\lesssim$ 0.1% of the total DM. We find that these constraints are competitive, and hence provide a new observable to probe the nature of PBH DM. The systematics affecting Lyman-$\alpha$ forest measurements are different from other constraining observations, and thus this is a complementary probe.
@@ I Introduction
Several astrophysical and cosmological evidences point towards the existence of Dark Matter (DM), something that makes up around $\sim$ 85% of matter content today. But all these evidences rely on the gravitational interactions of DM. Any non-gravitational nature of DM is yet to be discovered. Due to a wide variety of DM models and a huge range in the masses of possible DM candidates, there has been a plethora of research on the possibilities and methods of observing any imprints of non-gravitational DM interaction with Standard Model (SM) particles or with itself. [1, 2, 3, 4].
@@
Primordial Black Hole [5, 6, 7, 8, 9] (PBH) is one of the most popular and widely discussed DM candidates [10, 11, 12]. Various production mechanisms for PBH DM have been proposed in the literature and various aspects concerning their formation have been studied [13, 14, 15, 16, 17, 6, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Irrespective of its production mechanism, a PBH would emit Hawking radiation, the spectrum of which depends on the BH mass, charge, and spin [8, 41, 42, 43, 44, 45, 46]. Several terrestrial and space-based telescopes have been used to directly search for Hawking evaporation from PBH DM. These efforts have resulted in very stringent constraints on the abundance of low-mass ($\sim$ $10^{15}$ g - $5\times 10^{17}$ g ) PBH DM [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Besides these direct searches for Hawking evaporation from PBH DM, a number of techniques have also been explored to search for indirect imprints of this radiation. SM particles emitted by evaporating PBH DM can dump energy into their surroundings. Effects of this exotic energy injection have been searched using various cosmological observables like the Cosmic Microwave Background [60, 61, 62, 63] and the global 21 cm signal [64, 65, 66, 67, 68, 69, 70].
@@
Lyman-$\alpha$ ($\sim$1216 Å) forest, which is a series of absorption lines seen in the spectra of distant quasars, remains one of the most important tracers of the diffuse inter-galactic medium (IGM) [71, 72, 73, 74, 75, 76]. Besides being sensitive to the values of various cosmological parameters and the nature of reionizing sources [77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92], observations of these Lyman-$\alpha$ absorption features are sensitive to the thermal and ionization history of the IGM [71, 72, 73, 74, 93, 94, 95, 96, 97, 98]. With the current and promising progress in Lyman-$\alpha$ forest observations, it is important to investigate the potential of these observations to uncover Beyond the Standard Model (BSM) physics. Several works have already demonstrated the power of Lyman-$\alpha$ forest in searching for various BSM scenarios [99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113]. In this work, these observations are used for the first time to put limits on the abundance of low-mass evaporating PBH DM.
@@
This paper is structured as follows. Section (II) gives an overview of the energy injection into the IGM by evaporating PBHs. In section (III), we briefly describe the evolution of the IGM temperture and ionization history in the presence of Hawking evaporating PBHs. The first part of section (IV) is a brisk tour of the Lyman-$\alpha$ forest observations in the context of IGM temperature measurements. Then it lists the various datasets of temperature measurements using this technique that we make use of. We discuss our results in section (V) and then conclude in section (VI).
@@ II Energy injection by evaporating PBH DM
A PBH with a given mass and angular momentum will evaporate via Hawking radiation [8]. The temperature of an uncharged spinning PBH is,
@@
with $a_{*}$, $J$, and $M_{\rm PBH}$ being the reduced spin parameter, magnitude of angular momentum, and the mass of the PBH, respectively, and $G$ is the gravitational constant. For a non-spinning PBH ($a_{*}=0$), the above expression simplifies to,
@@
The number of $i^{\rm th}$ particle, $N_{i}$ emitted per unit energy per unit time by a Hawking evaporating BH is given by  [41, 42, 44, 45]
@@
where $\Gamma_{i}$ are the greybody factors [41, 42, 44] encoding the probability of an emitted particle overcoming the gravitational potential of the black hole so as to escape away from the black hole, and $E$ is the total energy of the emitted particle, $m$ is the axial quantum number, $s$ is the spin of the emitted particle, $\mu$ is the rest mass of the emitted particle, and dofdof stands for the degrees of freedom associated with the emitted particle.
@@
Figure 2: Constraints on non-spinning ($a_{*}$ = 0) (left panel) and spinning ($a_{*}$ = 0.9999) (right panel) PBH DM (monochromatic mass distribution). Constraints from the conservative model are shown in red (deep red and the shaded region above that), whereas the photoheating model limits are shown in blue (deep blue and the shaded region above that). The deep blue and deep red bands are obtained by varying over the photoheating models used in this work (FlexKnot [4], Tanh [4], and Trac [114, 115]). Previous constraints include low-energy positron measurements from Galactic Centre by INTEGRAL (navy [49] and green [118]), cosmic-ray flux measurement by Voyager (purple) [118], INTEGRAL (orange, dark red) and COMPTEL (teal) measurements of the Galactic gamma-ray flux [52, 59, 51], Super-Kamiokande limit for diffuse supernovae neutrino background (yellow) [53], measurement of extra-Galactic gamma-ray emission (aqua) [54, 55], gas heating in Leo T (black) [56, 57], PLANCK measurement of CMB (lime) [60], and XMM-NEWTON measurement of Galactic diffuse X-ray emission (pink) [118]. Note that two of the bounds - namely the cosmic-ray bounds from COMPTEL and INTEGRAL - for non-spinning PBH DM are not depicted for spinning PBH DM simply because the latter have not been derived in the literature.
In this work, the Hawking evaporation spectra are obtained using the publicly available code, BlackHawk v2.2[45, 119]. We have explicitly matched the spectra from BlackHawk with ref. [41]. At low energies, for computing secondary photon and electron-positron spectra, BlackHawk uses public Python code Hazma[120]. In this work we have used the total emission spectra (sum of primary and secondary) from PBHs in the mass range  $\sim\,5\times 10^{15}\,{\rm g}-10^{17}$ g. The lower limit on the mass of the PBHs comes from the fact that PBHs lighter than this mass lose a significant fraction of their mass in all the cosmological time from their formation epoch till today. Thus, their mass cannot be assumed to be constant, and thus, the emission spectra from them is time-dependent. We have probed those PBHs that have contributed the same to the dark matter density since their formation. These PBHs’ mass doesn’t change, and thus the evaporation spectrum of each emitted particle also remains the same throughout. Note that the only species of particles whose spectra we had to consider were photons and $e^{\pm}$. This is because only these particles are capable of efficiently interacting with the gas in the IGM and substantially affecting its thermal and ionization state [121, 122]. Other stable particles that are abundantly present in the Hawking radiation emitted by black holes of these masses are neutrinos and anti-neutrinos, and these have no effect on the temperature evolution of the IGM due to their weak-interaction.
@@
where $dV$ is the volume element and $f_{c}(z)$ is the deposition efficiency into channel $c$ at a given redshift zz[124, 125, 126], respectively. The relevant channels are hydrogen ionization ($f_{ion}$), helium ionization ($f_{He}$), hydrogen excitation ($f_{exc}$), and heating of IGM ($f_{heat}$). The $f_{c}(z)$’s have been computed in refs. [124, 125, 126] and have been implemented in the publicly available code DarkHistory[127, 106, 128, 129, 130, 131, 132]. In our work we have used DarkHistory to calculate the evolution of IGM temperature and ionization in presence of evaporating PBHs. Darkhistory enables us to calculate the the modified thermal and ionization state of the IGM self-consistently. What this means is that the effect of a modified ionization history at any redshift $z$ due to evaporating PBHs affects the heat deposition rates into the IGM also, and vice-versa. We have computed both the modified ionization history and the temperature history fully taking into account each other’s evolution (called backreaction in [106].)
@@
$\times\left(\frac{d^{2}N}{dEdt}\right)_{e^{\pm}}n_{\rm PBH}(M_{\rm PBH})\,\Psi(M_{\rm PBH})\,dM_{\rm PBH}\,dE_{e}.$ (9)
@@
with $f_{\rm PBH}$, $\rho_{c}$, and $\Omega_{\rm DM}$ being the fraction of DM in the form of PBHs, the critical density, and DM density parameter in the present universe. In our work we have assumed monochromatic mass distributions of PBH DM, i.e., all PBHs are assumed to be of the same mass, but the value of this mass is varied. Thus, $\Psi(M_{\rm PBH})$ takes the form of a dirac delta function peaked at a variable PBH mass $M_{\rm PBH}$.
III Ionization and Thermal history of the IGM
Having discussed the deposition of the injected energy sourced by Hawking evaporation of PBHs into the IGM, we now describe the effects of this deposited energy on the IGM temperature and the ionized hydrogen fraction. The rate of change of the IGM temperature $(\dot{T_{m}}$) is given by [133, 134, 135, 127],
@@ III Ionization and Thermal history of the IGM
In the above equations, $H$ is the Hubble parameter, $T_{\text{CMB}}$ is the Cosmic Microwave Background (CMB) temperature, $x_{e}$ is the free electron fraction ($\equiv n_{\text{e}}/{n_{\text{H}}}$). The parameter, $f_{\rm He}\equiv n_{\text{He}}/{n_{\text{H}}}$, where $n_{\text{He}}$ and $n_{\text{H}}$ are the number densities of helium and hydrogen nuclei, respectively; $\sigma_{T}$ is the Thomson cross section, $a_{r}$ is the radiation constant, and $m_{e}$ is the electron mass. The term $\dot{T}_{\text{atom}}$ in Eq. (12) represents the cooling term due to all atomic processes like collisional ionization, collisional excitation and Coulomb heating [124]. Finally, the $\dot{T}^{\star}$ term in Eq. (12) is the photoheating term that accounts for the heating due to photoionization from various astrophysical sources. This term is dictated by the ionization history of the IGM, which we briefly describe in what follows.
The evolution equation for the ionized hydrogen fractionin the IGM is:
@@
where $x_{\text{HII}}=n_{\text{HII}}/{n_{\text{H}}}$, with $n_{\text{HII}}$ being the number density of ionized hydrogen. The first term on the right hand side is again the contribution due to the standard astrophysical sources and the second term represents the contribution due to PBH energy injection. The first term in Eq. (15) is:
@@
Here, $x_{\text{e}}^{\text{P}}$ is taken from the Planck 95% confidence region limits on free electron fraction using ‘Tanh’ and ‘FlexKnot’ models [4]. In addition, we also use the 95% confidence region limit from parametrization described in Refs. [114, 115], hereafter referred to as the ‘Trac’ model. We note that, at sufficiently high redshifts  [^1] Please refer to Ref. [106] for more details on how to fix the value of redshift above which $\dot{x}_{\text{HII}}^{\star}=0$. $\dot{x}_{\text{HII}}^{\star}$ is taken to be zero., $\dot{x}_{\text{HII}}^{\star}=0$.
Finally, we use the following parametrization to fix $\dot{T}^{\star}$,
[…]
In the above equation, $\Delta T$ is a free parameter whose value is known to lie in the range $2\times 10^{4}\,{\rm K}-3\times 10^{4}$ K as discussed in Refs. [139, 140, 94, 141, 98]. We shall use the value of $\Delta T$ within this range which fits our temperature measurement data the best. Furthermore, $\nu_{\text{X}}$ is the frequency corresponding to the ionization potential of species $\text{X}\in\{\text{HI},\text{HeI}\}$, $h$ is Planck’s constant, $\gamma_{\text{X}}$ is the index for photoionization cross-section, $\alpha_{\text{bkg}}$ is the index for power-law ionizing background [140, 98, 97], and $\alpha_{\text{A},\text{X}}$ are the case-A recombination factors associated with the transitions XI $\rightarrow$ X [98]. Like $\Delta T$, $\alpha_{\rm bkg}$ is a free parameter which can be varied to get a best-fit to the temperature data, and whose value is also known to lie within a range- 0.5<1.5-0.5<1.5 in this case.
@@
where $\mathcal{R}=13.6$ eV is the hydrogen ionization potential and $\mathcal{C}$ is the Peebles’ C factor [133, 135].